# Quantum (quadratic) gravity: replacing the massive tensor ghost with an inverted harmonic oscillator-like instability

### 07 Mar 2026 v1

@@ 4.1.2 Optical theorem and unitarity
In momentum space, one introduces invariant amplitudes $\mathcal{M}_{fi}$ and the Lorentz-invariant phase-space measure $d\Pi_{f}$ so that (4.31) becomes
@@
Let us introduce the mathematical identities that would help us to frame the optical theorem in the language of the propagator and the on-shell contributions of the intermediate states. For a real variable $x$, the following distributional identity separates the pole at $x=0$ as
$\frac{1}{x-i\epsilon}=\mathrm{PV}\!\left(\frac{1}{x}\right)+i\pi\,\delta(x)\,,\qquad\epsilon\to 0^{+},$ (4.33)
where $\rm PV$ stands for the Principal Value, which can be understood as
$\mathrm{PV}\!\left(\frac{1}{x}\right)=\frac{1}{2}\left(\frac{1}{x-i\epsilon}+\frac{1}{x+i\epsilon}\right)={\rm Re}\left(\frac{1}{x-i\epsilon}\right)=\lim_{\varepsilon\to 0^{+}}\frac{x}{x^{2}+\varepsilon^{2}}\ \,$, (4.34)
which implies
$2\,\text{Im}\!\left(\frac{1}{x-i\epsilon}\right)=2\pi\,\delta(x).$ (4.35)
Applied to Feynman denominator of a massive scalar field,
@@
More generally, for any Feynman diagram $\mathcal{M}$ built from a Hermitian interaction Lagrangian, we find that the discontinuity $2i\,\text{Im}\left(\mathcal{M}\right)$ is given by a sum over cuts of the diagram, where each cut propagator is replaced by an on-shell factor proportional to $\delta(q^{2}+m^{2})\theta(q^{0})$. This is the Cutkosky cutting rule and is the diagrammatic implementation of inserting $\sum_{f}|f\rangle\langle f|$ in (4.31). Thus, for a given internal mode, whether it contributes to the optical theorem is determined by whether its on-shell delta-function has support on the physical kinematics.
@@ 4.1.3 The spin–2 ghost and unitarity violation
It is important to review how the optical theorem is violated with a ghost. Consider a forward elastic amplitude mediated (at one loop) by a pair of massive spin–2 modes, schematically as in Fig. 3. In a standard unitary theory, the discontinuity across the branch cut is obtained by putting the intermediate states on shell (Cutkosky rules), yielding
@@
$\frac{1}{k^{2}+m_{2}^{2}-i\epsilon}=\mathrm{PV}\!\left(\frac{1}{k^{2}+m^{2}}\right)+i\pi\,\delta(k^{2}+m_{2}^{2}),$ (4.40)
@@
To see unitarity violation, and how the minus sign emerges from the $\delta$-part (explicitly), in the most transparent way, it is enough to isolate the distributional (on-shell) part of a propagator. For any real mass $m_{2}$,
$\frac{1}{x\pm i\epsilon}=\mathrm{PV}\!\left(\frac{1}{x}\right)\mp i\pi\,\delta(x),\qquad x\equiv k^{2}+m_{2}^{2}.$ (4.44)
Hence a healthy Feynman propagator takes the form
$\Delta_{F}(k)\equiv\frac{-i}{k^{2}+m_{2}^{2}-i\epsilon}=-i\,\mathrm{PV}\!\left(\frac{1}{k^{2}+m_{2}^{2}}\right)+\pi\,\delta(k^{2}+m_{2}^{2}).$ (4.45)
The point is that the entire absorptive part (the part that contributes to $\mathrm{Im}$ of amplitudes) originates from the $\delta$-term. The principal value part is purely dispersive.
Now consider instead a ghost field, i.e. a mode whose quadratic action yields the opposite-sign propagator (negative residue). At the level relevant for unitarity, this is simply
$\Delta_{F}^{(g)}(p)\equiv\,\frac{i}{k^{2}+m_{2}^{2}-i\epsilon}=\,i\,\mathrm{PV}\!\left(\frac{1}{k^{2}+m_{2}^{2}}\right)-\pi\,\delta(k^{2}+m_{2}^{2}).$ (4.46)
Comparing (4.45) and (4.46), we see that the on-shell piece flips sign:
@@
This sign flip is the microscopic origin of the negative contribution to the optical theorem.
@@
$i\mathcal{M}^{\text{(bubble)}}(P)=\int\!\frac{d^{4}k}{(2\pi)^{4}}\;\big[iV_{L}\big]\;\Delta_{1}(k)\;\Delta_{2}(P-k)\;\big[iV_{R}\big],$ (4.48)
where $V_{L,R}$ are real tree-level vertices and $P$ is the external (forward) momentum. For healthy particles,
$\Delta_{1}(k)=\frac{-i}{k^{2}+m_{2}^{2}-i\epsilon},\qquad\Delta_{2}(P-k)=\frac{-i}{(P-k)^{2}+m_{2}^{2}-i\epsilon}.$ (4.49)
Using (4.45), the product of propagators contains a term
@@
which is real and positive as a distribution. This is precisely the term that yields the imaginary part of the loop amplitude after the $k^{0}$-integration: the two $\delta$-functions enforce on-shell conditions and produce the Lorentz-invariant phase space. More explicitly, the discontinuity (or twice the imaginary part) is obtained by the replacement
$\frac{-i}{k^{2}+m_{2}^{2}-i\epsilon}\;\longrightarrow\;2\pi\,\delta(k^{2}+m_{2}^{2})\qquad(\text{Cutkosky rule}),$ (4.51)
so that
$2\,\mathrm{Im}\,\mathcal{M}^{\text{(bubble)}}(P)=\int d\Pi_{2}(P)\;\Big|\mathcal{M}^{\text{(tree)}}_{i\to 2}(P)\Big|^{2},$ (4.52)
which is manifestly non-negative. For ghost lines, each propagator is multiplied by 1-1 since
$\Delta^{(g)}(p)=\frac{i}{k^{2}+m_{2}^{2}-i\epsilon}.$ (4.53)
Therefore the cut replacement (4.51) becomes
$\Delta^{(g)}(p)\;\longrightarrow\;-\,2\pi\,\delta(k^{2}+m_{2}^{2}),$ (4.54)
because the sign flip is entirely carried by the $\delta$-part as shown in (4.46)–(4.47). Cutting two ghost lines gives an overall factor $(-1)^{2}=+1$ from the propagators, but the crucial sign relative to the optical theorem comes from how the ghost couples to the Hilbert-space metric: the intermediate state sum effectively inserts an extra minus sign, equivalently traced to the negative residue/spectral weight. [^6] In operator language, the completeness relation carries the metric $\eta_{n}=\pm 1$ for an indefinite-norm state space, $\sum_{n}\eta_{n}|n\rangle\langle n|=1$. This is exactly the same sign that appears in the cut.
@@ 4.2 The saga of quantum prescriptions to save unitarity
One way to resolve ghost issue is to treat the massive spin–2 ghost in higher-derivative theories as a Lee–Wick–type resonance that becomes unstable (the idea of unstable ghost resonance) once loop effects are included [54, 51]. In that context, radiative corrections generate a width, the would be ghost is removed from the asymptotic spectrum, and perturbative unitarity can be organized by summing only over cuts of stable states; when one nevertheless works in a narrow-width approximation that treats the unstable mode as if it were cut, one must adopt the original Lee–Wick contour prescription to recover the correct result. However, a key technical obstacle to interpreting this as ordinary “decay” lies in the location of the dressed poles. For an ordinary unstable particle, the complex poles migrate to the second Riemann sheet (so the one-particle pole disappears from the physical sheet), whereas for a ghost resonance the dressed propagator generically develops a complex-conjugate pair of poles in the first (physical) sheet, which signals that the ghost cannot be eliminated as an asymptotic degree of freedom in the same way. This is precisely the point emphasized in recent analysis of Buoninfante [38] on ghost resonances where it was shown that above threshold, the ghost’s complex poles remain on the physical sheet (unlike standard resonances), leading to the conclusion that the naive “ghost decays away” picture is not technically justified in the usual analytic structure, and motivating the “no-decay/anti-instability” interpretation. A complementary viewpoint [76, 77] is the nonperturbative/IR-emergence perspective: rather than “fixing” the spin–2 ghost by a Lee–Wick contour or by treating it as a decaying resonance, it is argued that the perturbative pole structure is not the right guide to the physical spectrum once quadratic gravity runs strong. In this picture, motivated by a QCD analogy, the would-be ghost can be removed or reshaped by nonperturbative dynamics, while GR emerges as the effective IR description. A related classical/semiclassical theme is that the dangerous higher-derivative instability can be parametrically slow and can be damped in expanding backgrounds, so that the problematic mode need not appear as a physical asymptotic excitation even before invoking any explicit pole prescription.
@@ 5 Unitary quadratic quantum gravity with IHO-like spin–2 sector
Having shown that the standard quantization of quadratic gravity violates the optical theorem due to the massive spin–2 ghost, we now present an alternative unitary formulation based on the dual IHO spin–2 sector and its quantization scheme (See Sec . 3.1). Worth emphasizing that IHO physics is universal as it is prevalent in SM, also in other areas of physics, chemistry, and mathematics. Here, we see that it is highly relevant to quantum gravity. In this approach, the IHO-like instability arises from an extra spin–2 mode with spacelike-momentum support that does not appear in asymptotic states or unitarity cuts, yet remains consistent with virtual (purely quantum) processes. In this section, we demonstrate how this framework leads to a consistent realization of QQG, in which renormalizability and unitarity coexist and the additional spin–2 sector is understood as a dynamical instability rather than a fundamental inconsistency.
@@
The field $u_{s}$ is massless spin–2 field whereas the field $v_{s}$ is now a dual IHO spin–2 field, which is nothing but an infinite collection of dual IHO (See Table. 1). The graviton propagator (4.5) in this context becomes
@@
A key issue in assessing the physical viability of dual IHO spin–2 modes is whether they contaminate observable scattering processes and violate the optical theorem. We argue that this need not be the case since such modes are supported exclusively in the spacelike region of momentum space. Spacelike excitations do not define asymptotic particle states and therefore cannot appear as external legs in the $S$-matrix. Their contributions are restricted to internal lines and are kinematically separated from the timelike sector describing ordinary matter and radiation. This separation allows the theory to maintain unitarity and causal propagation for physical (on-shell) degrees of freedom, while accommodating spacelike modes (dual IHO spin–2) as effective degrees of freedom associated with instabilities, horizons, or IHO dynamics.
5.1 The dual IHO spin–2 field, unitarity cuts and renormalization
In particular, the spin–2 sector has the schematic form
$D^{(2)}_{\mu\nu,\rho\sigma}(k)=-\frac{i}{M_{P}^{2}}\,P^{(2)}_{\mu\nu,\rho\sigma}\left[\frac{1}{k^{2}-i\epsilon}-\frac{1}{k^{2}+m_{2}^{2}-i\epsilon}\right],$ (5.5)
where $m_{2}^{2}$ is the additional spin–2 mass parameter determined by the quadratic coupling $\beta$. In the case of IHO spin–2 case, we have
$m_{2}^{2}=-\mu_{2}^{2}<0\qquad\Rightarrow\qquad k^{2}+m_{2}^{2}=k^{2}-\mu_{2}^{2},$ (5.6)
so the additional spin–2 pole is located at the spacelike surface $k^{2}=\mu_{2}^{2}>0$. The dual-IHO spin–2 propagator can be factorized as
$\frac{1}{k^{2}-\mu_{2}^{2}-i\epsilon}=\mathrm{PV}\!\left(\frac{1}{k^{2}-\mu_{2}^{2}}\right)+i\pi\,\delta\!\left(k^{2}-\mu_{2}^{2}\right)\,$, (5.7)
we see explicitly that only the $\delta(k^{2}-\mu_{2}^{2})$ term can contribute to the absorptive part (i.e. to unitarity cuts involved in the optical theorem sum). The principal value piece is purely real and therefore contributes only to the real part of amplitudes.
In the following, we show that the dual IHO spin–2 mode is not subject to the optical theorem. In quadratic gravity with a spin–2 dual IHO pole at $k^{2}=\mu_{2}^{2}>0$, it does not contribute to the optical theorem for physical graviton scattering: it cannot appear as an on-shell intermediate state in any unitarity cut. By the cutting identity (4.35), an internal line with denominator $(k^{2}-\mu_{2}^{2}-i\epsilon)^{-1}$ can contribute to $2\,\text{Im}\left(\mathcal{M}\right)$ only through the on-shell distribution
$2\,\text{Im}\!\left(\frac{1}{k^{2}-\mu_{2}^{2}-i\epsilon}\right)=2\pi\,\delta(k^{2}-\mu_{2}^{2}).$ (5.8)
Thus, the IHO mode can enter the optical theorem only if the physical kinematics allow $k^{2}=\mu_{2}^{2}$ for some cut momentum $k$. For $2\to 2$ a (forward) elastic scattering of physical massless external gravitons with momenta $p_{1},p_{2}\to p_{3},p_{4}$ obeying $p_{i}^{2}=0$ and $p_{i}^{0}>0$, one has the kinematic inequalities
$s\leq 0,\qquad t\geq 0,\qquad u\geq 0,$ (5.9)
with strict $s<0$ in the center-of-mass frame (since $k_{s}=(E_{1}+E_{2},\mathbf{0})$ implies $s=k_{s}^{2}=-(E_{1}+E_{2})^{2}$). Although the momentum transfers $k_{t},k_{u}$ are spacelike ($t,u\geq 0$), the optical theorem for forward $2\to 2$ scattering is controlled by the cut momentum flowing through the intermediate state, which is the total incoming momentum $k=k_{s}=p_{1}+p_{2}$. Therefore the on-shell condition for the dual-IHO pole, $k^{2}=\mu_{2}^{2}>0$, cannot be satisfied in the physical forward region, and equivalently
$\delta(k_{s}^{2}-\mu_{2}^{2})=\delta(s-\mu_{2}^{2})=0\qquad(\mu_{2}^{2}>0)\,.$ (5.10)
Therefore, the dual IHO spin–2 propagator yields no contribution to $2\,\mathrm{Im}(M)$ in the forward unitarity cut. Nevertheless, this spin–2 mode can still contribute off shell to tree amplitudes and to loop renormalization (real parts and UV divergences) through the principal value pieces of propagators, but it cannot contribute to the unitarity (optical theorem) sum because it never goes on shell in physical kinematics. It is important to note that, at the tree level, the dual-IHO exchange contributes only through the principal value part in (5.7), hence only to the real (dispersive) part of the amplitude. There is no absorptive part associated with putting the IHO mode on shell, and hence no IHO spin–2 contribution to the optical theorem.
We have, therefore, a sharp separation between (i) dispersive/off-shell effects of the dual-IHO spin–2 pole and (ii) absorptive/on-shell contributions that enter the optical theorem. This is illustrated in Fig. 4.
(i) The dual IHO spin–2 field and the off-shell effects.
$i\Pi(P)\;=\;\int\!\frac{d^{4}k}{(2\pi)^{4}}\,\frac{\mathcal{N}(k,P)}{\big(k^{2}-\mu^{2}-i\epsilon\big)\big((P-k)^{2}-\mu^{2}-i\epsilon\big)}.$ (5.11)
The location of branch points and cuts is controlled entirely by the Landau (pinch) conditions, i.e. by the possibility of putting the internal lines simultaneously on shell. For the present bubble these are
$k^{2}=\mu_{2}^{2},\qquad(P-k)^{2}=\mu_{2}^{2}.$ (5.12)
Subtracting the two equations in (5.12) gives
$k^{\mu}=a\,P^{\mu}+k_{\perp}^{\mu},\qquad P\!\cdot\!k_{\perp}=0.$ (5.14)
Then (5.13) fixes $a=\tfrac{1}{2}$, i.e.
$k^{\mu}=\frac{1}{2}P^{\mu}+k_{\perp}^{\mu},\qquad P\!\cdot\!k_{\perp}=0.$ (5.15)
Imposing the on-shell condition $k^{2}=\mu^{2}$ and using (5.15) yields
$\mu^{2}=k^{2}=\left(\frac{P}{2}\right)^{2}+k_{\perp}^{2}=\frac{P^{2}}{4}+k_{\perp}^{2}\quad\Longrightarrow\quad k_{\perp}^{2}=\mu_{2}^{2}-\frac{P^{2}}{4}.$ (5.16)
$P^{2}_{\ast}=4\mu_{2}^{2}\,$}. (5.17)
(ii) Renormalization without extra unitarity channels and principal value prescription.
In loop integrals the internal momentum $k$ is integrated over all values, so the factor $\delta(k^{2}-\mu_{2}^{2})$ in (5.7) can only contribute if it appears in a unitarity cut of a diagram, i.e. if the loop discontinuity is supported on a pinch singularity where the relevant internal lines can be simultaneously put on shell with the appropriate $\theta(k^{0})$ restrictions. The previous discussion states that this cannot occur for physical graviton scattering since $k^{2}=\mu_{2}^{2}>0$ is kinematically inaccessible for the channel momenta that control the physical discontinuities.
For any computations of loop effects in the QQG, the dual-IHO prescription consists of replacing the propagator with the principal value since the delta function part drops out of the unitarity cuts for forward amplitudes. Using (4.33) with $x=k^{2}-\mu_{2}^{2}$ gives
$\frac{1}{k^{2}-\mu_{2}^{2}-i\epsilon}=\mathrm{PV}\!\left(\frac{1}{k^{2}-\mu_{2}^{2}}\right)+i\pi\,\delta(k^{2}-\mu_{2}^{2}).$ (5.18)
The dual-IHO (principal value) prescription consists of retaining only the principal value part and removing the $\delta$-supported on-shell contribution:
$\frac{1}{k^{2}-\mu_{2}^{2}-i\epsilon}\;\longrightarrow\;\mathrm{PV}\!\left(\frac{1}{k^{2}-\mu_{2}^{2}}\right)\,.$ (5.19)
Thus, the spin–2 propagator used in internal lines becomes
$D^{(2)}_{\mu\nu\rho\sigma}(k)=\frac{i}{M_{p}^{2}}\,P^{(2)}_{\mu\nu\rho\sigma}\left[\frac{1}{k^{2}-i\epsilon}-\mathrm{PV}\!\left(\frac{1}{k^{2}-\mu_{2}^{2}}\right)\right]\,$, (5.20)
which has no $\delta(k^{2}-\mu_{2}^{2})$ support from the extra spin–2 pole.
$\mathcal{M}_{1\text{-loop}}(p_{i})\;\sim\;\int\!\frac{d^{4}q}{(2\pi)^{4}}\;\frac{N(q,p_{i})}{\prod_{a}\big(D_{a}(q,p_{i})-i\epsilon\big)}.$ (5.21)
Here $p_{i}$ are external momenta, and:
  1. 1.
  2. 2.
To isolate the effect of a single IHO spin–2 line, it is convenient to rewrite (5.21) as
$\mathcal{M}_{1\text{-loop}}\;\sim\;\int\!\frac{d^{4}q}{(2\pi)^{4}}\;F(q;p_{i})\;\frac{1}{q^{2}-\mu_{2}^{2}-i\epsilon},$ (5.22)
where we define
$F(q;p_{i})\equiv\frac{N(q,p_{i})}{\prod_{a\neq\text{IHO}}\big(D_{a}(q,p_{i})-i\epsilon\big)}\,.$ (5.23)
In words: $F(q;p_{i})$ is simply all the remaining in the integrand after factoring out the IHO-like spin–2 in the denominator. For fixed external kinematics, $F(q;p_{i})$ is a (typically smooth) function of $q$ away from the poles of the other denominators. Now, we proceed to identify the exact principal value and $\delta$ decomposition at the level of the loop. Applying (4.33) to (5.22) yields the exact splitting
$\mathcal{M}_{1\text{-loop}}\;\sim\;\int\!\frac{d^{4}q}{(2\pi)^{4}}\;F(q;p_{i})\;\mathrm{PV}\!\left(\frac{1}{q^{2}-\mu_{2}^{2}}\right)\;+\;i\pi\int\!\frac{d^{4}q}{(2\pi)^{4}}\;F(q;p_{i})\;\delta(q^{2}-\mu_{2}^{2}).$ (5.24)
The second term is the on-shell contribution associated with the $\delta$-support. Under the dual-IHO prescription (5.19), one drops precisely this term:
$\mathcal{M}_{1\text{-loop}}^{\text{(dual-IHO)}}\;\sim\;\int\!\frac{d^{4}k}{(2\pi)^{4}}\;F(k;p_{i})\;\mathrm{PV}\!\left(\frac{1}{k^{2}-\mu_{2}^{2}}\right)\,.$ (5.25)
Let us now verify that, in the UV limit, the principal-value prescription preserves the renormalizability of quadratic gravity. Using (4.34), for $|k^{2}|\gg\mu_{2}^{2}$ one finds
Therefore, at large loop momentum, the dual IHO line falls off as $1/k^{2}$ (up to subleading powers), just like an ordinary propagator. Consequently, the UV degree of divergence of (5.25) is controlled by the same power counting as in the standard Feynman prescription; the IHO mode still contributes to UV divergences, counterterms, and running couplings.
which is supported only on the (finite) hypersurface $k^{2}=\mu_{2}^{2}$ in momentum space. This on-shell term generates the discontinuities (absorptive contributions) associated with cuts through that line, but it does not control the $|k|\to\infty$ region that produces UV divergences. Hence the dual-IHO prescription removes the on-shell (cut-producing) part of the dual IHO line while retaining its off-shell contribution to renormalization and real (dispersive) loop corrections. Consequently, the dual-IHO mode can and does contribute virtually through the principal value part of its propagator. This yields:
  1. 1.
  2. 2.
    $D^{(2)}(k)\;\propto\;\left(\frac{1}{k^{2}-i\epsilon}-\frac{1}{k^{2}-\mu_{2}^{2}-i\epsilon}\right)=\frac{\mu_{2}^{2}}{(k^{2}-i\epsilon)(k^{2}-\mu_{2}^{2}-i\epsilon)}\,$, (5.28)
In summary, the dual-IHO prescription enforces a clean split between what is allowed and what is forbidden. The IHO-like spin–2 part is off-shell as there is no $\delta$-function support, thus its propagator is understood as a principal value, so loop integrals generate only real contributions, exactly the pieces needed for renormalization (counterterms and the real parts of self-energies/vertices). Since the $\delta$-supported part is what normally produces an absorptive contribution and hence a unitarity cut (via the optical theorem), the corresponding physical channel involving the spin–2 IHO never opens. The dual-IHO spin–2 mode is therefore present as a virtual, renormalizing sector but is excluded from the spectrum of on-shell intermediate states.
[…]
$D^{(2)}(k)$
(b) UV improvement via $1/k^{4}$ falloff.
Figure 4: Dual-IHO spin–2: absent from physical unitarity cuts (top), but improves UV behavior via $1/q^{4}$ falloff (bottom).

### 25 Mar 2026 v2

@@ 4.1.2 Optical theorem and unitarity
where
$\langle f|T|i\rangle=(2\pi)^{4}\delta^{4}\!\left(\sum_{j\in\mathrm{final}}p_{j}-\sum_{i\in\mathrm{initial}}p_{i}\right)\cdot\mathcal{M}_{fi}$ (4.32)
Expressing the optical theorem (4.31) in terms of the invariant amplitudes $\mathcal{M}_{fi}$ and the Lorentz-invariant phase-space measure $d\Pi_{f}$, we get
@@
For the optical theorem, an important point is that the imaginary part of the invariant amplitude gets a contribution from the delta function part of the Feynman propagator (here we assume a case of a massive scalar field) via
@@
where $F(q)$ is a generic function of the momenta that depends on the nature of the loops and vertices in a particular physical process.
For any Feynman diagram $\mathcal{M}$ built from a Hermitian interaction Lagrangian, we find that the discontinuity $2\,\text{Im}\left(\mathcal{M}\right)$ is given by a sum over cuts of the diagram, where each cut propagator is replaced by an on-shell factor proportional to $\delta(q^{2}+m^{2})\theta(q^{0})$. This is the Cutkosky cutting rule and is the diagrammatic implementation of inserting $\sum_{f}|f\rangle\langle f|$ in (4.31). Thus, for a given internal mode, whether it contributes to the optical theorem is determined by whether its on-shell delta-function has support on the physical kinematics.
@@ 4.1.3 The spin–2 ghost and unitarity violation
It is important to review how the optical theorem is violated with a ghost.
Tree-level unitarity violation:
Let us consider the case of 2-2 (massless) graviton scattering with massive (ghost) spin–2. The invariant amplitude in this case becomes
$i\mathcal{M}=\left(-iV^{\mu\nu\rho\sigma}(k_{1},k_{2})\right)\times\frac{i\,P^{(2)}_{\mu\nu\alpha\beta}}{M_{p}^{2}(k^{2}+m_{2}^{2}-i\epsilon)}\times\left(-iV^{\alpha\beta}_{\rho\sigma}(k_{3},k_{4})\right)$ (4.37)
where $V^{\mu\nu\rho\sigma}$ are vertex factors comes from the 3rd order perturbation of the quadratic gravity action (4.1) around Minkowski $g_{\mu\nu}=\eta_{\mu\nu}+\kappa h_{\mu\nu}$. Simplifying further (4.37) gives
$\mathcal{M}=-\frac{\,\mathcal{N}}{M_{p}^{2}(k^{2}+m_{2}^{2}-i\epsilon)}=\frac{\mathcal{N}}{M_{p}^{2}}\left[\mathrm{PV}\left(\frac{1}{k^{2}+m_{2}^{2}}\right)+i\pi\delta(k^{2}+m_{2}^{2})\right]$ (4.38)
where $\mathcal{N}=V^{\mu\nu\rho\sigma}P^{(2)}_{\mu\nu\alpha\beta}V^{\alpha\beta}_{\rho\sigma}>0$ collects all real positive kinematic factors from the vertices and spin-2 projector contracted on physical external graviton polarizations. Let us now verify the optical theorem (4.33) below
$\underbrace{2\,\mathrm{Im}(\mathcal{M}^{\mathrm{tree,ghost}}_{gg\to gg})}_{\mathrm{LHS}}=-\frac{2\pi\,\mathcal{N}}{M_{p}^{2}}\delta(k^{2}+m_{2}^{2})\neq+\frac{2\pi\,\mathcal{N}}{M_{p}^{2}}\delta(k^{2}+m_{2}^{2})=\underbrace{\int d\Pi_{1}\,|\mathcal{M}^{\mathrm{tree}}_{gg\to v}|^{2}}_{\mathrm{RHS}}$ (4.39)
from which we can conclude that the optical theorem is clearly violated.
Loop-level unitarity violation:
Consider a forward elastic amplitude mediated (at one loop) by a pair of massive spin–2 modes, schematically as in Fig. 3. In a standard unitary theory, the discontinuity across the branch cut is obtained by putting the intermediate states on shell (Cutkosky rules), yielding
@@
$\frac{1}{k^{2}+m_{2}^{2}-i\epsilon}=\mathrm{PV}\!\left(\frac{1}{k^{2}+m_{2}^{2}}\right)+i\pi\,\delta(k^{2}+m_{2}^{2}),$ (4.41)
@@
To see unitarity violation, and how the minus sign emerges from the $\delta$-part (explicitly), in the most transparent way, it is enough to isolate the distributional (on-shell) part of a propagator. In other words, the following sign flip is the microscopic origin of the negative contribution to the optical theorem.
@@
$i\mathcal{M}^{\text{(bubble)}}(P)=\int\!\frac{d^{4}k}{(2\pi)^{4}}\;\big[-iV_{L}\big]\;\Delta_{1}(k)\;\Delta_{2}(P-k)\;\big[-iV_{R}\big],$ (4.46)
where $V_{L,R}$ are real tree-level vertices and $P$ is the external (forward) momentum. For the ghost (spin–2) particles,
$\Delta_{1}(k)=\frac{i}{k^{2}+m_{2}^{2}-i\epsilon},\qquad\Delta_{2}(P-k)=\frac{i}{(P-k)^{2}+m_{2}^{2}-i\epsilon}.$ (4.47)
The absorptive contributions come from product of propagators contains a term
@@
which is real and positive as a distribution. This is precisely the term that yields the imaginary part of the loop amplitude after the $k^{0}$-integration: the two $\delta$-functions enforce on-shell conditions and produce the Lorentz-invariant phase space.
Cutting two ghost lines gives an overall factor $(-1)^{2}=+1$ from the propagators, but the crucial sign relative to the optical theorem comes from how the ghost couples to the Hilbert-space metric: the intermediate state sum effectively inserts an extra minus sign, equivalently traced to the negative residue/spectral weight. [^6] In operator language, the completeness relation carries the metric $\eta_{n}=\pm 1$ for an indefinite-norm state space, $\sum_{n}\eta_{n}|n\rangle\langle n|=1$. This is exactly the same sign that appears in the cut.
@@ 4.2 The saga of quantum prescriptions to save unitarity
One way to resolve ghost issue is to treat the massive spin–2 ghost in higher-derivative theories as a Lee–Wick–type resonance that becomes unstable (the idea of unstable ghost resonance) once loop effects are included [56, 53]. In that context, radiative corrections generate a width, the would be ghost is removed from the asymptotic spectrum, and perturbative unitarity can be organized by summing only over cuts of stable states; when one nevertheless works in a narrow-width approximation that treats the unstable mode as if it were cut, one must adopt the original Lee–Wick contour prescription to recover the correct result. However, a key technical obstacle to interpreting this as ordinary “decay” lies in the location of the dressed poles. For an ordinary unstable particle, the complex poles migrate to the second Riemann sheet (so the one-particle pole disappears from the physical sheet), whereas for a ghost resonance the dressed propagator generically develops a complex-conjugate pair of poles in the first (physical) sheet, which signals that the ghost cannot be eliminated as an asymptotic degree of freedom in the same way. This is precisely the point emphasized in recent analysis [40] on ghost resonances where it was shown that above threshold, the ghost’s complex poles remain on the physical sheet (unlike standard resonances), leading to the conclusion that the naive “ghost decays away” picture is not technically justified in the usual analytic structure, and motivating the “no-decay/anti-instability” interpretation [96, 97]. A complementary viewpoint [78, 79] is the nonperturbative/IR-emergence perspective: rather than “fixing” the spin–2 ghost by a Lee–Wick contour or by treating it as a decaying resonance, it is argued that the perturbative pole structure is not the right guide to the physical spectrum once quadratic gravity runs strong. In this picture, motivated by a QCD analogy, the would-be ghost can be removed or reshaped by nonperturbative dynamics, while GR emerges as the effective IR description. A related classical/semiclassical theme is that the dangerous higher-derivative instability can be parametrically slow and can be damped in expanding backgrounds, so that the problematic mode need not appear as a physical asymptotic excitation even before invoking any explicit pole prescription.
@@ 5 Unitary quadratic quantum gravity with IHO-like spin–2 sector
Having shown that the standard quantization of quadratic gravity violates the optical theorem due to the massive spin–2 ghost, we now present an alternative unitary formulation based on the dual-IHO spin–2 sector and its quantization scheme (See Sec . 3.1). Worth emphasizing that IHO physics is universal as it is prevalent in SM, also in other areas of physics, chemistry, and mathematics. Here, we see that it is highly relevant to quantum gravity. In this approach, the IHO-like instability arises from an extra spin–2 mode with spacelike-momentum support that does not appear in asymptotic states or unitarity cuts, yet remains consistent with virtual (purely quantum) processes. In this section, we demonstrate how this framework leads to a consistent realization of QQG, in which renormalizability and unitarity coexist and the additional spin–2 sector is understood as a dynamical instability rather than a fundamental inconsistency.
@@
The field $u_{s}$ is massless spin–2 field whereas the field $v_{s}$ is now a dual-IHO spin–2 field, which is nothing but an infinite collection of dual-IHO (See Table. 1). The graviton propagator (4.5) in this context becomes
@@
A key issue in assessing the physical viability of dual-IHO spin–2 modes is whether they contaminate observable scattering processes and violate the optical theorem. We argue that this need not be the case since such modes are supported exclusively in the spacelike region of momentum space. Spacelike excitations do not define asymptotic particle states and therefore cannot appear as external legs in the $S$-matrix. Their contributions are restricted to internal lines and are kinematically separated from the timelike sector describing ordinary matter and radiation. This separation allows the theory to maintain unitarity and causal propagation for physical (on-shell) degrees of freedom, while accommodating spacelike modes (dual-IHO spin–2) as effective degrees of freedom associated with instabilities, horizons, or IHO dynamics.